# Neutrino observables in gauged $U(1)_{L_{\alpha}-L_{\beta}}$ models with two Higgs doublet and one singlet scalars

### v1

@@ I Introduction
An introduction of new $U(1)$ gauge symmetry provides simple extensions of the SM. Among many possibilities local $U(1)_{L_{\alpha}-L_{\beta}}$ symmetries are interesting where they are anomaly free and the SM leptons have flavor dependent charge $L_{\alpha}-L_{\beta}(\alpha,\beta=\{e,\mu,\tau\})$. Introduction of such a symmetry restricts the structure of Yukawa interactions and Majorana mass of right-handed neutrinos that are related to neutrino mass structure. Actually compatibility among the $U(1)_{L_{\alpha}-L_{\beta}}$ symmetry and neutrino masses/mixings has been explored for type-I seesaw scenario Asai:2018ocx ; Asai:2017ryy where one scalar field, $SU(2)_{L}$ singlet or doublet with non-zero $U(1)_{L_{\alpha}-L_{\beta}}$ charge, is introduced to develop a vacuum expectation value (VEVs) contributing to neutrino mass via Yukawa interaction as minimal choices  [^1] One can find other approaches, especially $U(1)_{L_{\mu}-L_{\tau}}$ case, discussing constrained neutrino mass matrix in other seesaw models including radiative mass generation Baek:2015mna ; Lee:2017ekw ; Chen:2017gvf ; Nomura:2018vfz ; Nomura:2018cle ; Araki:2019rmw ; Asai:2019ciz ; Bauer:2020itv ; Nomura:2023vmh . . In this case we have clear relations among neutrino observables and models are restricted by the current neutrino data Esteban:2020cvm . It is then found that only $U(1)_{L_{\mu}-L_{\tau}}$ case with one singlet scalar having charge ±$\pm 1$ can accommodate neutrino data when we consider constraints on sum of neutrino mass $\sum m_{\nu}\lesssim 0.12$ eV from CMB data by Planck Planck:2018vyg under the standard $\Lambda$CDM cosmological model. Remarkably more stringent constraint is obtained as $\sum m_{\nu}<0.07$ eV (95%) if we include recent baryon acoustic oscillation (BAO) analysis by Dark Energy Spectroscopic Instrument (DESI) data DESI:2024mwx ; even stronger bounds are estimated in refs. Craig:2024tky ; Wang:2024hen . Then $U(1)_{L_{\alpha}-L_{\beta}}$ models are excluded when we only add one new scalar field and impose the constraint on the sum; it is marginal even if we consider $\sum m_{\nu}>0.059$ eV prior leading bit looser constraint $\sum m_{\nu}<0.113$. These analysis indicate we need to extend minimal $U(1)_{L_{\alpha}-L_{\beta}}$ models or modify standard $\Lambda$CDM model. Thus it is worth considering a next minimal case including both $SU(2)_{L}$ singlet and doublet scalar fields  [^2] Such a hybrid type is discussed in ref. Asai:2024pzx to relax constraints regarding $Z^{\prime}$ boson. and investigate if we can still get some predictions in neutrino sector satisfying the constraint on neutrino mass at the same time.
@@ II.2 Neutral gauge bosons
In this work we are not going to discuss $Z^{\prime}$ physics assuming $Z$-$Z^{\prime}$ mixing is negligibly small to avoid electroweak precision tests, and require $Z^{\prime}$ mass and gauge coupling $g_{X}$ satisfying current constraints from $Z^{\prime}$ searches.
@@ II.3 Charged lepton mass
where $\times$ indicates a non-zero element. The mass matrix can be diagonalized by bi-unitary matrices $V_{eL}$ and $V_{eR}$ transforming $\ell_{L}$ and $\ell_{R}$; $M_{d}^{\rm diagonal}=V_{eL}M_{e}V_{eR}^{\dagger}$. Note that such transformation induces FCNC for interaction among $Z^{\prime}$ and leptons due to lepton flavor dependent charges and it is strongly constrained by LFV search experiments Asai:2018ocx . We then assume off-diagonal elements of the matrix is negligibly small and $V_{eL(eR)}\simeq 1$ for simplicity.
@@ II.4 Neutrino mass
After spontaneous symmetry breaking we obtain Dirac mass $M_{D}$ between $\nu_{L}$ and $\nu_{R}$ as well as Majorana mass $M_{R}$ of $\nu_{R}$. These mass matrices are written by
@@ III Numerical analysis
Fig. 3 and 4 show $\delta_{\rm CP}$-$\alpha_{21}$ and $\delta_{\rm CP}$-$\alpha_{31}$ values for allowed parameter points in each model. We find some correlations among phases where correlations of $\delta_{\rm CP}$-$\alpha_{21}$ in models (1) and (6) [(3) and (5)] provide similar behavior. In each case the value of $\alpha_{21}$ is concentrated around 100-250 degrees. On the other hand correlations of $\delta_{\rm CP}$-$\alpha_{31}$ are different in each model. In these plots we show 1(3)$\sigma$ range of $\delta_{\rm CP}$ value by the light-blue(cyan) region as in the Fig. 2. In addition, we also show correlation between $\alpha_{21}$ and $\alpha_{31}$ in Fig. 5 where models (1) and (6) [(3) and (5)] provide similar behavior.

### v2

@@ I Introduction
An introduction of new $U(1)$ gauge symmetry provides simple extensions of the SM. Among many possibilities local $U(1)_{L_{\alpha}-L_{\beta}}$ symmetries are interesting where they are anomaly free and the SM leptons have flavor dependent charge $L_{\alpha}-L_{\beta}(\alpha,\beta=\{e,\mu,\tau\})$. Introduction of such a symmetry restricts the structure of Yukawa interactions and Majorana mass of right-handed neutrinos that are related to neutrino mass structure. Actually compatibility among the $U(1)_{L_{\alpha}-L_{\beta}}$ symmetry and neutrino masses/mixings has been explored for type-I seesaw scenario Asai:2018ocx ; Asai:2017ryy where one scalar field, $SU(2)_{L}$ singlet or doublet with non-zero $U(1)_{L_{\alpha}-L_{\beta}}$ charge, is introduced to develop a vacuum expectation value (VEVs) contributing to neutrino mass via Yukawa interaction as minimal choices  [^1] One can find other approaches, especially $U(1)_{L_{\mu}-L_{\tau}}$ case, discussing constrained neutrino mass matrix in other seesaw models including radiative mass generation Baek:2015mna ; Lee:2017ekw ; Chen:2017gvf ; Nomura:2018vfz ; Nomura:2018cle ; Araki:2019rmw ; Asai:2019ciz ; Bauer:2020itv ; Nomura:2023vmh . . In this case we have clear relations among neutrino observables and models are restricted by the current neutrino data Esteban:2020cvm . It is then found that only $U(1)_{L_{\mu}-L_{\tau}}$ case with one singlet scalar having charge ±$\pm 1$ can accommodate neutrino data when we consider constraints on sum of neutrino mass $\sum m_{\nu}\lesssim 0.12$ eV from CMB data by Planck Planck:2018vyg under the standard $\Lambda$CDM cosmological model. Remarkably more stringent constraint is obtained as $\sum m_{\nu}<0.07$ eV (95%) if we include recent baryon acoustic oscillation (BAO) analysis by Dark Energy Spectroscopic Instrument (DESI) data DESI:2024mwx ; even stronger bounds are estimated in refs. Craig:2024tky ; Wang:2024hen . Then $U(1)_{L_{\alpha}-L_{\beta}}$ models are excluded when we only add one new scalar field and impose the constraint on the sum; it is marginal even if we consider $\sum m_{\nu}>0.059$ eV prior leading bit looser constraint $\sum m_{\nu}<0.113$. These analysis indicate we need to extend minimal $U(1)_{L_{\alpha}-L_{\beta}}$ models or modify standard $\Lambda$CDM model. Thus it is worth considering a next minimal case including both $SU(2)_{L}$ singlet and doublet scalar fields  [^2] Such a hybrid type is discussed in ref. Asai:2024pzx to relax constraints regarding $Z^{\prime}$ boson. Also similar hybrid models are discussed in refs. Heeck:2011wj ; Heeck:2014qea ; Crivellin:2015lwa ; Ardu:2022zom for lepton flavor violation physics. and investigate if we can still get some predictions in neutrino sector satisfying the constraint on neutrino mass at the same time.
@@ II.2 Neutral gauge bosons
Here we briefly discuss constraints on $Z$-$Z^{\prime}$ mixing from electroweak precision test (EWPT). The mixing is typically constrained as $|\sin\chi|\lesssim 10^{-3}-10^{-4}$ when $Z^{\prime}$ boson comes from local $U(1)$ symmetry as subgroup of GUT symmetry like $SO(10)$ and $E_{6}$ Erler:2009jh . EWPT constraint is discussed for $U(1)_{L_{\mu}-L_{\tau}}$ case in ref. Heeck:2011wj where we find $-0.0008\lesssim g_{X}\sin\chi\lesssim 0.0003$. We apply the constraint as a reference under $m_{Z}^{2}\ll m_{Z^{\prime}}^{2}$ mass hierarchy for neutral gauge bosons. In our case the constraint approximately becomes
$\frac{g_{X}^{2}m_{Z}v\cos\beta}{m_{Z^{\prime}}^{2}}\lesssim 8\times 10^{-4}.$ (17)
In Fig. 1, we show the region excluded by EWPT for $U(1)_{L_{\mu}-L_{\tau}}$ with $\tan\beta=5$. It is shown that $\mathcal{O}(1)$ gauge coupling is allowed if $Z^{\prime}$ mass satisfies 1700 GeV $\lesssim m_{Z^{\prime}}$ where the allowed region is also safe from other experimental constraints. We expect similar $Z$-$Z^{\prime}$ mixing constraint in the other cases, $U(1)_{L_{e}-L_{\mu}}$ and $U(1)_{L_{e}-L_{\tau}}$. This constraint restricts $v_{\varphi}\simeq m_{Z^{\prime}}/g_{X}$ and can affect Majorana mass of $\nu_{R}$. However it does not provide much impact in fitting neutrino mass since we also have Yukawa couplings that are free parameters. In our analysis below we consider heavy $Z^{\prime}$ and assume $Z$-$Z^{\prime}$ mixing is negligibly small to avoid EWPT.
Figure 1: The gray region is excluded by EWPT for $U(1)_{L_{\mu}-L_{\tau}}$ case with $\tan\beta=5$.
@@ II.3 Charged lepton mass
where $\times$ indicates a non-zero element. The mass matrix can be diagonalized by bi-unitary matrices $V_{eL}$ and $V_{eR}$ transforming $\ell_{L}$ and $\ell_{R}$; $M_{d}^{\rm diagonal}=V_{eL}^{\dagger}M_{e}V_{eR}$. Note that such transformation induces FCNC for interaction among $Z^{\prime}$ and leptons due to lepton flavor dependent charges and it is strongly constrained by LFV search experiments Asai:2018ocx . Here we illustrate a LFV constraint in a simple case. Firstly $Z^{\prime}$ interaction for charged leptons is written by
$\mathcal{L}_{\rm Z^{\prime}\ell\bar{\ell}}=g_{X}Z^{\prime}_{\mu}\left[\overline{\ell_{L}}V^{\dagger}_{eL}Q_{L_{\alpha}-L_{\beta}}V_{eL}\gamma^{\mu}\ell_{L}+\overline{\ell_{R}}V^{\dagger}_{eR}Q_{L_{\alpha}-L_{\beta}}V_{eR}\gamma^{\mu}\ell_{R}\right],$ (39)
where $Q_{L_{\alpha-\beta}}$ indicates diagonal matrix whose elements are $L_{\alpha}-L_{\beta}$ charge, e.g. $Q_{L_{e}-L_{\mu}}=(1,-1,0)$ diagonal matrix. The FCNC appears when $V^{\dagger}_{eL(eR)}Q_{L_{\alpha}-L_{\beta}}V_{eL(eR)}$ is not diagonal matrix. As an example we consider $M_{e}$ of model (3) in Eq. (38) with $(M_{e})_{23}=0$ for simplicity. Then, focusing on $e$-$\mu$ sector, the mass matrix is written by
$M_{e}^{e\mu}=\begin{pmatrix}M_{e}&\delta m\\ 0&M_{\mu}\end{pmatrix}.$ (40)
Choosing $\delta m\ll M_{\mu}$ this matrix can be approximately diagonalized by
$V^{\dagger}_{eL}M_{e}^{e\mu}V_{eR}\sim\begin{pmatrix}m_{e}&0\\ 0&m_{\mu}\end{pmatrix},\quad V_{eL}\sim\begin{pmatrix}1&\frac{\delta m}{M_{\mu}}\\ -\frac{\delta m}{M_{\mu}}&1\end{pmatrix},\quad V_{eR}\sim\bf{1},$ (41)
where $m_{e}\sim M_{e}$ and $m_{\mu}\sim M_{\mu}$. Since model (3) is $U(1)_{L_{e}-L_{\tau}}$ case the LFV interaction is approximately given by
$g_{X}\frac{\delta m}{m_{\mu}}Z^{\prime}_{\mu}[\bar{\mu}\gamma^{\mu}P_{L}e+\bar{e}\gamma^{\mu}P_{L}\mu].$ (42)
Then the experimental constraint for $\mu\to ee\bar{e}$ process is most stringent one for heavy $Z^{\prime}$ case  [^3] The case of light $Z^{\prime}$ allowing $\ell_{i}\to\ell_{j}Z^{\prime}$ decay is discussed in ref. Asai:2018ocx providing tighter constraints.; that is induced via virtual $Z^{\prime}$ as $\mu\to eZ^{\prime*}\to ee\bar{e}$. We can estimate the decay width and branching ratio (BR) as
$\Gamma_{\mu\to ee\bar{e}}\simeq\frac{g_{X}^{4}}{8m_{Z^{\prime}}^{4}}\left(\frac{\delta m}{m_{\mu}}\right)^{2}\frac{m_{\mu}^{5}}{192\pi^{3}},$ (43)
$BR(\mu\to ee\bar{e})\simeq\frac{g_{X}^{4}}{g_{2}^{2}G_{F}^{2}m_{Z^{\prime}}^{4}}\left(\frac{\delta m}{m_{\mu}}\right)^{2},$ (44)
where $G_{F}$ is the Fermi constant. Applying the current limit $BR(\mu\to ee\bar{e})\lesssim 10^{-12}$ SINDRUM:1987nra , we obtain
$g_{X}^{4}\left(\frac{\delta m}{m_{\mu}}\right)^{2}\left(\frac{\rm TeV}{m_{Z^{\prime}}}\right)^{4}\lesssim 6\times 10^{-11}.$ (45)
Thus $\delta m/m_{\mu}$ should be very small as $\delta m/m_{\mu}\lesssim\mathcal{O}(10^{-5})$ when $g_{X}=\mathcal{O}(1)$ and $m_{Z^{\prime}}$ is TeV scale. Note that we will obtain looser constraint when we consider LFV decay of $\tau$ lepton. In the numerical analysis below we assume off-diagonal elements of the matrix is negligibly small and $V_{eL(eR)}\simeq 1$ for simplicity.
@@ II.4 Neutrino mass
After spontaneous symmetry breaking we obtain Dirac mass $M_{D}$ between $\nu_{L}$ and $\nu_{R}$ ($\overline{\nu_{L}}M_{D}\nu_{R}$) as well as Majorana mass $M_{R}$ of $\nu_{R}$. These mass matrices are written by
@@ III Numerical analysis
Fig. 4 and 5 show $\delta_{\rm CP}$-$\alpha_{21}$ and $\delta_{\rm CP}$-$\alpha_{31}$ values for allowed parameter points in each model. We find some correlations among phases where correlations of $\delta_{\rm CP}$-$\alpha_{21}$ in models (1) and (6) [(3) and (5)] provide similar behavior. In each case the value of $\alpha_{21}$ is concentrated around 100-250 degrees. On the other hand correlations of $\delta_{\rm CP}$-$\alpha_{31}$ are different in each model. In these plots we show 1(3)$\sigma$ range of $\delta_{\rm CP}$ value by the light-blue(cyan) region as in the Fig. 3. In addition, we also show correlation between $\alpha_{21}$ and $\alpha_{31}$ in Fig. 6 where models (1) and (6) [(3) and (5)] provide similar behavior.